Joan planned a workout program that included increasing the amount of time she exercised by 1/5 hour each week for the first 5 weeks
if Joan starts Week 1 with a 1/4 hour workout how long will her workout program be at the beginning of week 5

Answer by ikleyn(52942)

.

Week 1: 1/4 of an hours, or 15 minutes.


Week 5: add the increment of 1/5 of an hour (which is 12 minutes) 4 times to get


        15 minutes + 4*12 minutes = 15 + 48 minutes = 63 minutes = 1 hour and 3 minutes.    ANSWER

Solved.


PLEASE  use the punctuation signs  (the commas)  in your posts to make them   R E A D A B L E.